OpenStudy (anonymous):

Please help i will reward a medal , Last 3 questions. :)

OpenStudy (anonymous):

Here to help what are they?

OpenStudy (anonymous):

Alright wait a sec

OpenStudy (anonymous):

y=-4x+6

OpenStudy (anonymous):

Yeah that right

OpenStudy (anonymous):

second one is 3/5

OpenStudy (anonymous):

okay i understand what i did wrong with that one i just checked

OpenStudy (anonymous):

ok i had that answer for the second one

OpenStudy (anonymous):

it's \[\frac{ 16 }{ 7 }x+y=-8\]

OpenStudy (anonymous):

No y=-16/7x-8

OpenStudy (anonymous):

They asked for standard form

OpenStudy (anonymous):

What never mind the other answer is right

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

Sorry guys my laptop crashed , thank you :) I really appreciate it.